OpenStudy (anonymous):

Determine whether the set S is linearly independent ir linearly depentdent. S = {(4,-3,6,2)(1,8,3,1)(3,-2,-1,0)}.

OpenStudy (anonymous):

this is an equivalent problem: is the following system consistent? : 4x + y + 3z = 0 -3x + 8y -2z = 0 6x + 3y - z =0 2x + y = 0

OpenStudy (amistre64):

row reduce the matrix

OpenStudy (anonymous):

or that

OpenStudy (anonymous):

@amistre64 is there a determinant method if we have n equations and n-1 unknowns ?

OpenStudy (amistre64):

there is but im not sure how to read it with nonsquare matrixes

OpenStudy (amistre64):

in essesne it comes across as a cross product instead of a singular value

OpenStudy (amistre64):

rref, and if you obtain any columns without a pivot entry, your linearlyD